OpenStudy (1018):

a conical container has a height of 20cm and a base diameter of 12cm. It holds water a depth of 10cm above its vertex. Find the volume of its content.

OpenStudy (anonymous):

Ok I got this, volume for conical container is V = 1/3 pi r2 h V=1/3(pi)(6^2)(20) Volume approximately equals 753.98

OpenStudy (anonymous):

753.98cm

OpenStudy (1018):

wait, can I ask why did you use the 20cm?

OpenStudy (1018):

and the answer's not in the choices. :(

OpenStudy (anonymous):

because that was the height you gave

OpenStudy (1018):

but it also gave me 10cm. what is that for?

OpenStudy (1018):

i tried that with the same givens before but gave me the wrong answer too :(

OpenStudy (anonymous):

is 251.32 one of your answers?

OpenStudy (1018):

that's a no too

OpenStudy (wolf1728):

I think you need the formula for the volume of a frustum Here's the formula: Frustum Volume = (π • height • (Ra2 + Ra • Rb + Rb2) ÷ 3) and here's a link to my web page: http://www.1728.org/volcone.htm

OpenStudy (wolf1728):

In order to use that formula we have to calculate the radius of the circle at 10 cm height Base radius = 6 height = 20 cm

OpenStudy (1018):

wait should i use 10 as height or 20?

OpenStudy (wolf1728):

I was busy making a graphic. We need to calculate the length of the red line.

OpenStudy (wolf1728):

Is the calculation as easy as taking one half of 6? The length of the red line is 3 ?

OpenStudy (wolf1728):

I get volume of 659.73 is that one of the choices?

OpenStudy (1018):

no :( haha

OpenStudy (1018):

i tried to use 3 and 10 in the volume formula and i got the answer. lol

OpenStudy (1018):

but im not sure if its right

OpenStudy (1018):

but its in the choices

OpenStudy (wolf1728):

no? wow I thought that was it. I used 3 and 6 for the radii and height of 10

OpenStudy (1018):

i thought so too. haha. do you know any other ways to get the upper radius?

OpenStudy (wolf1728):

If you don't mind, what are the answers?

OpenStudy (1018):

the choices are 20 30 40 60, all with pi (e.g 20pi)

OpenStudy (wolf1728):

The biggest answer is 60*PI which is about 188.5 which isn't even half of the answer I gave. Is the problem stated corrrectly? Height 20 Halfway 10 base radius =6 top radius = 3

OpenStudy (1018):

yup that's correct. but we dont know the top radius if its really 3

OpenStudy (wolf1728):

I'd say it is 3

OpenStudy (1018):

what if the conical section is upside down? so the content would be from the tip and only one radius? haha

OpenStudy (1018):

that's what i though when i used 3 and 10 lol

OpenStudy (1018):

It holds water a depth of 10cm above its vertex does this means something else?

OpenStudy (wolf1728):

I always thought the vertex was the absolute top point of a cone , triangle, etc. Maybe I'll see what wikipedia says.

OpenStudy (1018):

yup that's right but you know, just some outside thinking. haha

OpenStudy (wolf1728):

Yes, it's what I said - the absolute top point of a cone.

OpenStudy (wolf1728):

By the way, I just tried 2 calculations to fine the volume volume of a 20 cm hgt 6cm radius cone 753.98 volume of a 10 cm hgt 3cm radius cone 94.248 Subtracting 1 from the other = 659.732 Hmm that looks familiar!!! LOL

OpenStudy (1018):

haha! are we close to the answer? LOL

OpenStudy (wolf1728):

no I don't think we are any closer. But at least you know HOW to find the volume.

OpenStudy (wolf1728):

By the way, why did you pick 1018 as a user name?

OpenStudy (1018):

wait, what do you mean. i know how to get volume. lol just random numbers. to stay anonymous. haha. i like your though. because wolf. nice.

OpenStudy (wolf1728):

Thanks - and the 1728 (besides the name of my website) is also 12 cubed = 12^3 = 12*12*12 1,728

OpenStudy (1018):

nice nice! mathematical. hahaha

OpenStudy (wolf1728):

Absolutely!!! By the way, feel free to use my website all you want. It has a lot more than just cone calculators. If you are taking algebra geometry rig, etc you;ll find all calculators to help you.

OpenStudy (wolf1728):

that "rig " should be trig LOL

OpenStudy (1018):

haha hey thanks!

OpenStudy (wolf1728):

Okay, you are welcome. Well, think I should go. Good luck with those math problems.

OpenStudy (wolf1728):

Thanks for the medal!!

OpenStudy (1018):

no problem!

OpenStudy (wolf1728):

Okay see ya!!